OpenStudy (anonymous):

Help with finding principal value. I've memorized the unit circle , but my teacher was very vague over principal values.

OpenStudy (anonymous):

i dont get how \[\sin^{-1} (-\sqrt{2}/2) is -\pi/4\]

OpenStudy (anonymous):

because \[\sin(-\frac{\pi}{4})=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

you are looking for the angle (number) in the interval \[[-\frac{\pi}{2},\frac{\pi}{2}]\] where the sine is your number

OpenStudy (anonymous):

there are an infinite number of numbers whose sine is anything between -1 and 1 because it is periodic, but only one in the above mentioned interval

OpenStudy (anonymous):

So it sin inverse would only be in the first quadrant of the unit circle?

OpenStudy (mertsj):

No. It could also be in the 4th quadrant. All values for sin^-1 are, by definition, in quadrants I or IV